Targaryen Corporation has a target capital structure of 70 percent common stock, 5 percent preferred stock, and 25 percent debt. Its cost of equity is 10 percent, the cost of preferred stock is 5 percent, and the pretax cost of debt is 6 percent. The relevant tax rate is 23 percent.

a.

What is the company's WACC? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the aftertax cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

aftertax cost of debt=6(1-tax rate)

=6(1-0.23)=4.62%

WACC=Respective costs*Respective weights

=(0.7*10)+(0.05*5)+(0.25*4.62)

=8.41%(Approx).